OpenStudy (anonymous):

The circle shown below has AB and BC as its tangents. If the measure of arc AC is 160°, what is the measure of angle ABC? 80° 40° 60° 20°

OpenStudy (anonymous):

Can you post the image please?

OpenStudy (anonymous):

can anyone help me with this question ?

OpenStudy (cwrw238):

the angle subtended at center by arc AC = 160 the kite formed as 1 angle = 160 and 2 = 90 degrees so angle ABC = 369 - (90+ 90 + 160)

OpenStudy (cwrw238):

can you work that out?

OpenStudy (cwrw238):

are you there?

OpenStudy (anonymous):

sorry i was verifying my email

OpenStudy (anonymous):

529?

OpenStudy (cwrw238):

thats alright

OpenStudy (cwrw238):

work out whats in the parentheses first: 360 -(90+90+160) = 360 - 340 = 20

OpenStudy (anonymous):

so the answer is 20?

OpenStudy (cwrw238):